The mass of Earth is approximately 6.0 x 1024 kg. What would the acceleration due to gravity be on the surface of planet Q which has a mass of 5.0 x 1024 kg and a radius equal to the radius of Earth?  

  


1
Expert's answer
2021-04-26T16:48:29-0400

Acceleration due to gravity is "G \frac{M}{r^2}".

For Earth, "g = G \frac{M_E}{r_E^2}", and for planet Q it is: "g' = G \frac{M_Q}{r_E^2} = g \cdot \frac{M_Q}{M_E} = 9.81 \frac{m}{s^2} \cdot \frac{5 \cdot 10^{24} kg}{6 \cdot 10^{24} kg} = 9.81 \cdot {5}{6} \approx 8.18 \frac{m}{s^2}".